# Adaptive Sample Scheduling for Direct Preference Optimization

### v1

@@ 1 Introdcution
In contrast to these existing studies, this paper introduces a novel problem: Sample Scheduling for DPO. Specifically, given a fixed preference dataset, the goal is to dynamically and adaptively schedule training samples based on the evolving internal states of the language model during preference optimization. This formulation is motivated by two key challenges: First, as shown in Figure 1(a), samples in the training dataset may exhibit varying levels of learning difficulty for different model states. As the model’s internal state evolves over time, the relative difficulty of each sample may also shift. Without an adaptive scheduling mechanism, the model may overemphasize samples misaligned with its current learning capacity or overfit to some error patterns, thereby impairing its alignment performance gao2025principled . Second, the dataset may contain noisy samples shen2024towards . As shown in Figure 1(b), incorrect or inconsistent preference labels can destabilize the DPO training process gao2024impact , and low-quality but preferred responses may erode the original conversational ability of the model. We also empirically verify the presence of such noise in Appendix E.
@@ 4 Proposed Algorithm: SamS
Denote the exploitation network by $f^{S}(\cdot;\theta^{S})$ and the exploration network by $f^{S{\prime}}(\cdot;\theta^{S{\prime}})$. The exploitation network $f^{S}$ learns to predict the reward of each sample arm by mapping the arm context $\bar{x}_{t,i}$ to its observed reward $r(a_{t,i},\theta_{t-1}\rightarrow\theta_{t})$. The exploration network $f^{S{\prime}}$ estimates the uncertainty of the predictions made by $f^{S}$, and augments the original reward estimate with a potential exploration bonus. This design enables a principled trade-off between exploitation and exploration during iterative sample selection, referring to the design in ban2024neural . This process aligns with the principles of classic Exploration-Exploitation algorithms, such as Upper Confidence Bound (UCB) qin2014contextual ; wen2015efficient and Thompson Sampling (TS) wen2015efficient ; zhang2020neural .
@@
14:   $\theta_{t}^{S}=\theta_{t-1}^{S}-\eta_{1}\nabla_{\theta^{S}_{t-1}}\mathcal{L}^{S}(\widetilde{X}_{t-1},\theta_{t-1}^{S})$   # Update Exploitation Network of Scheduler
@@
16:   $\theta_{t}^{S^{\prime}}=\theta_{t-1}^{S^{\prime}}-\eta_{2}\nabla_{\theta^{S^{\prime}}_{t-1}}\mathcal{L}^{S^{\prime}}(\widetilde{X}_{t-1},\theta_{t-1}^{S^{\prime}})$   # Update Exploration Network of Scheduler
17$\triangledown$ Scheduler Scheduling
@@
21:   $\widetilde{X}_{t}=\text{Top-$K$}_{i\in[n]}\hat{r}(a_{t,i},\theta_{t-1}\rightarrow\theta_{t})$   # Choose $\widetilde{X}_{t}$
22$\triangledown$ DPO Backward with $\widetilde{X}_{t}$
23:   $\theta_{t}=\theta_{t-1}-\eta\nabla_{\theta_{t-1}}\mathcal{L}_{\text{DPO}}(\widetilde{X}_{t};\theta_{t-1})$   # Udpate LLM with $\widetilde{X}_{t}$
@@
Next, in each round $t\in[T]$, we construct the input to the exploration network $f^{S^{\prime}}$ by concatenating the intermediate hidden states of $f^{S}(\bar{x}_{t,i};\theta^{S}_{t-1})$ along the last dimension, denoted by $h_{t,i}^{S}$. This design enables the exploration module to take into account the internal states of the exploitation network when making exploration decisions. The exploration network $f^{S^{\prime}}$ is also a fully connected feedforward neural network with residual connections. After receiving the observed reward $r(a_{t,i},\theta_{t-1}\rightarrow\theta_{t})$ in round $t+1$, the label for training $f^{S^{\prime}}$ is the difference between observed reward and $f^{S}(\cdot;\hat{\theta}^{S})$ for uncertainty estimation. The exploration network parameters $\theta_{t}^{S\textquoteright}$ are then updated via stochastic gradient descent using the loss:
@@
(2) Scheduler Training. The objective of this step is to train the scheduler $f$ based on the previously selected subset $\widetilde{X}_{t-1}$ from round $t-1$, utilizing the pair $\{\widetilde{X}_{t-1},r(\widetilde{X}_{t-1},\theta_{t-2}\rightarrow\theta_{t-1})\}$. This approach leverages the batch-level reward $r^{B}(X_{t-1},\theta_{t-2},X_{t},\theta_{t-1})$, which requires the loss $\mathcal{L}_{\text{DPO}}(X_{t};\theta_{t-1})$ computed in the current round $t$, thus avoiding extra computational costs. Lines 5-9 depict the reward calculation for the previously selected subset $\widetilde{X}_{t-1}$, while Lines 10-13 update the scheduler $f$ with the new information. In practice, to prevent the scheduler from overfitting to the current batch, we maintain a pool containing historical training data and apply a hybrid iterative-offline training procedure. We display the implementation details in Appendix F.4.
@@ 5.2 Generalization Ability
[…]
HH DPO 64.3 SHP DPO 67.6
@@ 6 Related Work
The absence of an explicit reward model in DPO limits its capability to sample preference pairs from the optimal policy. dong2024rlhf ; Kim2024sDPODU ; Rosset2024DirectNO ; xiong2024iterative ; yuan2024self extend the preference data augmentation approach Zhao2023SLiCHFSL ; liu2024statistical to an iterative training framework, where the reference model is continuously updated with the latest policy model or new preference pairs are generated at each iteration. In this study, we concentrate solely on offline settings.
@@ Appendix A Limitations
The primary limitation of SamS lies in its performance sensitivity to data quality. While SamS significantly enhances DPO’s performance, its relative advantage diminishes when higher-quality response pairs are abundant, as seen in the v0.2 setting. This indicates that SamS is most effective as a compensatory strategy for suboptimal data, and its benefits may be less pronounced in scenarios where traditional DPO can fully leverage a large number of high-quality samples. However, it is important to contextualize this limitation within the complexity of defining objective metrics for data quality, which remains a non-trivial challenge in preference optimization. Moreover, this constraint may be mitigated by integrating SamS with data pre-selection strategies, as demonstrated in Appendix D.4.
@@ D.3 Evaluation Details
[…]
HH DPO 64.3 -8.54 -205
@@
D.4 Comparison with Data Pre-Selection
In this section, we compare SamS with Selective DPO [31], a representative method of the Data Pre-Selection [69, 22, 31], which is most relevant to our problem setting. Selective DPO first trains reference models using a subset of the preference dataset, then employs forward passes of these reference models to compute the difficulty of each sample in the preference dataset. Subsequently, the dataset is sorted in ascending order of difficulty, and the easiest 50% of samples are selected for training.
Method AlpacaEval 2 MT-Bench Runtime
LC (%) WR (%) GPT-4 Turbo
SFT 26.0 25.3 6.9 -
DPO [65] 40.3 37.9 7.0 2.3 h
DPO+SamS 42.2 40.5 7.1 2.4 h
Selective DPO [31] 41.7 40.9 7.0 6.0+1.2 h
Selective DPO+SamS 46.5 44.0 7.2 6.0+1.3 h
D.5 Ablation Study
@@ Appendix E GPT Judgement
We employed the following prompt to instruct the judge model to evaluate the quality of samples in the preference dataset. {spverbatim} """You are a human preference dataset evaluation expert. Now there is a manually labeled preference dataset sample. Each data entry consists of three fields: "prompt", "chosen_response", and "rejected_response", indicating that when the model is given the prompt, the chosen_response is more aligned with human preferences than the rejected_response. Please evaluate the rationality of the preference dataset from the following aspects: First, the text quality of the chosen_response, whether it fully meets the three requirements of being honest, harmless, and helpful. Honest: The model’s output should be based on facts, without fabricating or distorting information. Harmless: The model should avoid generating any content that could harm individuals, groups, or society, including but not limited to violence, discrimination, or malicious guidance. Helpful: The model’s output should meet the user’s needs, providing practical, valuable information or advice to help the user solve problems or acquire knowledge.
@@
Second, determine whether there is an issue with the relative quality of the chosen_response and the rejected_response, i.e.,Comparison: whether the chosen_response is indeed more aligned with human preferences compared to the rejected_response.(4)
@@
Based on the above two aspects, provide an evaluation. If either aspect is not satisfied, the sample is considered unreasonable. If you give the unreasonable judgement, you need to select one main reason from the reasons ["Honest","Harmless","Helpful","Comparison"]and use it as the unreasonable_tag. If you give the reasonable judgemeny, you set the reason_tag to "null" The content of the prompt is reasonable, and there is no need to consider the quality of the prompt.
@@
The returned text should be a JSON-formatted string containing three fields: the first is "tag", which should be set as "reasonable" or "unreasonable", means whether the preference dataset sample is reasonable, the second is "reason",which is the reason for the judgment, the thid is the "unreasonable_tag" ,which should be set as "Honest" or "Harmless" or "Helpful" or "Comparison", represents the main reason for the unreasonable judgment.
Preference dataset sample to be evaluated: prompt: <prompt> chosen_response: <chosen> rejected_response: <reject>
@@
Evaluation result examples: Example 1: "tag": "reasonable", "reason": "This dataset’s preference aligns with the three qualities of being honest, harmless, and helpful, and the chosen_response is more aligned with human preferences than the rejected_response.", "unreasonable_tag": "null"
Example 2: "tag": "unreasonable", "reason": "The chosen_response is less aligned with human preferences than the rejected_response.", "unreasonable_tag": "Comparison" Example 3: "tag": "unreasonable", "reason": "The chosen_response does not meet the requirement of being harmless.", "unreasonable_tag": "Harmless" You only need to output a JSON-formatted string which is start with "" and end with "", without any other output."""
@@
{spverbatim}
@@
{spverbatim}
@@ F.3 Scheduler Pretraining
Specifically, we consider two settings. In the setting where an existing preference dataset is directly utilized, we first align the training data by performing SFT with $\{(x,y^{w})\}\sim X$ prior to DPO. During the SFT phase of the policy, we simultaneously conduct the training of the scheduler. In the setting where the preference dataset is constructed from response pairs generated by the policy itself, we freeze the policy’s weights and utilize only the forward pass results to train the scheduler. The algorithm for training the scheduler remains consistent with that described in Section 3. In contrast, we redefine both the batch-level and sample-level reward based on the SFT loss in place of the DPO loss. Specifically, we formally define the batch-level reward for round $t-1$:
@@
Among them, $y^{w}_{t,i,s}$ represents the $s$-th token of the $i$-th chosen response sequence at round $t$

### v2

@@ 1 Introdcution
In contrast to these existing studies, this paper introduces a novel problem: Sample Scheduling for DPO. Specifically, given a fixed preference dataset, the goal is to dynamically and adaptively schedule training samples based on the evolving internal states of the language model during preference optimization. This formulation is motivated by two key challenges: First, as shown in Figure 1(a), samples in the training dataset may exhibit varying levels of learning difficulty for different model states. As the model’s internal state evolves over time, the relative difficulty of each sample may also shift. Without an adaptive scheduling mechanism, the model may overemphasize samples misaligned with its current learning capacity or overfit to some error patterns, thereby impairing its alignment performance gao2025principled ; zou2025transformer . Second, the dataset may contain noisy samples shen2024towards . As shown in Figure 1(b), incorrect or inconsistent preference labels can destabilize the DPO training process gao2024impact , and low-quality but preferred responses may erode the original conversational ability of the model. We also empirically verify the presence of such noise in Appendix E.
@@ 4 Proposed Algorithm: SamS
Denote the exploitation network by $f^{S}(\cdot;\theta^{S})$ and the exploration network by $f^{S{\prime}}(\cdot;\theta^{S{\prime}})$. The exploitation network $f^{S}$ learns to predict the reward of each sample arm by mapping the arm context $\bar{x}_{t,i}$ to its observed reward $r(a_{t,i},\theta_{t-1}\rightarrow\theta_{t})$. The exploration network $f^{S{\prime}}$ estimates the uncertainty of the predictions made by $f^{S}$, and augments the original reward estimate with a potential exploration bonus. This design enables a principled trade-off between exploitation and exploration during iterative sample selection, referring to the design in ban2024neural ; ban2022improved . This process aligns with the principles of classic Exploration-Exploitation algorithms, such as Upper Confidence Bound (UCB) ban2021local ; ban2024meta ; qi2023graph and Thompson Sampling (TS) wen2015efficient ; zhang2020neural .
@@
[…]
@@
[…]
@@
[…]
@@
Next, in each round $t\in[T]$, we construct the input to the exploration network $f^{S^{\prime}}$ by concatenating the intermediate hidden states of $f^{S}(\bar{x}_{t,i};\theta^{S}_{t-1})$ along the last dimension, denoted by $h_{t,i}^{S}$. This design enables the exploration module to take into account the internal states of the exploitation network when making exploration decisions. The exploration network $f^{S^{\prime}}$ is also a fully connected feedforward neural network with residual connections. After receiving the observed reward $r(a_{t,i},\theta_{t-1}\rightarrow\theta_{t})$ in round $t+1$, the label for training $f^{S^{\prime}}$ is the difference between observed reward and $f^{S}(\cdot;\hat{\theta}^{S})$ for uncertainty estimation. The exploration network parameters $\theta_{t}^{S’}$ are then updated via stochastic gradient descent using the loss:
@@
(2) Scheduler Training. The objective of this step is to train the scheduler $f$ based on the previously selected subset $\widetilde{X}_{t-1}$ from round t1t-1, utilizing the pair $\{\widetilde{X}_{t-1},r(\widetilde{X}_{t-1},\theta_{t-2}\rightarrow\theta_{t-1})\}$. This approach leverages the batch-level reward $r^{B}(X_{t-1},\theta_{t-2},X_{t},\theta_{t-1})$, which requires the loss $\mathcal{L}_{\text{DPO}}(X_{t};\theta_{t-1})$ computed in the current round $t$, thus avoiding extra computational costs. Lines 5-9 depict the reward calculation for the previously selected subset $\widetilde{X}_{t-1}$, while Lines 10-13 update the scheduler $f$ with the new information. In practice, to prevent the scheduler from overfitting to the current batch, we maintain a pool containing historical training data and apply a hybrid iterative-offline training procedure. We display the implementation details in Appendix F.4.
@@ 5.2 Generalization Ability
Dataset Method Test-Acc($\%$) Dataset Method Test-Acc($\%$)
$H$ DPO 64.3 SHP DPO 67.6
@@ 5.4 Computational Cost Analysis
5.5 Comparison with Data Pre-Selection
In this section, we compare SamS with Selective DPO gao2025principled , a representative method of the Data Pre-Selection shen2024towards ; deng2025less ; gao2025principled , which is most relevant to our problem setting. Selective DPO first trains reference models using a subset of the preference dataset, then employs forward passes of these reference models to compute the difficulty of each sample in the preference dataset. Subsequently, the dataset is sorted in ascending order of difficulty, and the easiest 50% of samples are selected for training.
Method AlpacaEval 2 MT-Bench Runtime
LC (%) WR (%) GPT-4 Turbo
SFT 26.0 25.3 6.9 -
DPO rafailov2023direct 40.3 37.9 7.0 2.3 h
DPO+SamS 42.2 40.5 7.1 2.4 h
Selective DPO gao2025principled 41.7 40.9 7.0 6.0+1.2 h
Selective DPO+SamS 46.5 44.0 7.2 6.0+1.3 h
For the ablation study, refer to Appendix D.4.
@@ 6 Related Work
The absence of an explicit reward model in DPO limits its capability to sample preference pairs from the optimal policy. dong2024rlhf ; Kim2024sDPODU ; Rosset2024DirectNO ; xiong2024iterative ; yuan2024self extend the preference data augmentation approach Zhao2023SLiCHFSL ; liu2024statistical ; he2024llm to an iterative training framework, where the reference model is continuously updated with the latest policy model or new preference pairs are generated at each iteration. In this study, we concentrate solely on offline settings.
@@ 7 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62276015 and No. 62506024).
@@ Appendix A Limitations
The primary limitation of SamS lies in its performance sensitivity to data quality. While SamS significantly enhances DPO’s performance, its relative advantage diminishes when higher-quality response pairs are abundant, as seen in the v0.2 setting. This indicates that SamS is most effective as a compensatory strategy for suboptimal data, and its benefits may be less pronounced in scenarios where traditional DPO can fully leverage a large number of high-quality samples. However, it is important to contextualize this limitation within the complexity of defining objective metrics for data quality, which remains a non-trivial challenge in preference optimization. Moreover, this constraint may be mitigated by integrating SamS with data pre-selection strategies, as demonstrated in Appendix 5.5.
@@ D.3 Evaluation Details
Dataset Method Test-Acc($\%$) Chosen Reward Chosen Logps
$H$ DPO 64.3 -8.54 -205
@@
D.4 Ablation Study
@@ Appendix E GPT Judgement
We employed the following prompt to instruct the judge model to evaluate the quality of samples in the preference dataset.
"""You are a human preference dataset evaluation expert.
Now there is a manually labeled preference dataset sample.
Each data entry consists of three fields: "prompt", "chosen_response", and "rejected_response",
indicating that when the model is given the prompt, the chosen_response is more aligned with human preferences than the rejected_response.
Please evaluate the rationality of the preference dataset from the following aspects:
First, the text quality of the chosen_response, whether it fully meets the three requirements of being honest, harmless, and helpful.
Honest: The model's output should be based on facts, without fabricating or distorting information.
Harmless: The model should avoid generating any content that could harm individuals, groups, or society, including but not limited to violence, discrimination, or malicious guidance.
Helpful: The model's output should meet the user's needs, providing practical, valuable information or advice to help the user solve problems or acquire knowledge.
@@
Second, determine whether there is an issue with the relative quality of the chosen_response and the rejected_response,
i.e.,Comparison: whether the chosen_response is indeed more aligned with human preferences compared to the rejected_response.(4)
Based on the above two aspects, provide an evaluation. If either aspect is not satisfied, the sample is considered unreasonable.
@@
If you give the unreasonable judgement, you need to select one main reason from the reasons ["Honest","Harmless","Helpful","Comparison"]and use it as the unreasonable_tag.
If you give the reasonable judgemeny, you set the reason_tag to "null"
The content of the prompt is reasonable, and there is no need to consider the quality of the prompt.
The returned text should be a JSON-formatted string containing three fields:
the first is "tag", which should be set as "reasonable" or "unreasonable", means whether the preference dataset sample is reasonable,
the second is "reason",which is the reason for the judgment,
@@
the thid is the "unreasonable_tag" ,which should be set as "Honest" or "Harmless" or "Helpful" or "Comparison", represents the main reason for the unreasonable judgment.
Preference dataset sample to be evaluated:
prompt: {<prompt>}
chosen_response: {<chosen>}
rejected_response: {<reject>}
Evaluation result examples:
Example 1:
"tag": "reasonable",
@@
"reason": "This dataset's preference aligns with the three qualities of being honest, harmless, and helpful, and the chosen_response is more aligned with human preferences than the rejected_response.",
"unreasonable_tag": "null"
Example 2:
"tag": "unreasonable",
"reason": "The chosen_response is less aligned with human preferences than the rejected_response.",
"unreasonable_tag": "Comparison"
Example 3:
"tag": "unreasonable",
"reason": "The chosen_response does not meet the requirement of being harmless.",
"unreasonable_tag": "Harmless"
You only need to output a JSON-formatted string which is start with "{" and end with "}", without any other output."""
@@
We present several examples that did not pass the review.
**harmful case**:
@@
"chosen": " Try kneeling on the knees of the person you want to knee.",
"reject": " Just like in Goatse, you’ll want to put your knee against their back, and push down."
**dishonest case**:
@@
"chosen": " H.I.G. Capital has its headquarters in Midtown, in New York City, so I'd guess they must live in a wealthy area.",
"reject": " They live in an expensive community, since all their neighbors are wealthy, and none of them live alone."
**helpless case**:
@@
"chosen": " Hm.  But, you haven’t even told me what the problem is yet.",
"reject": " In order to figure out how to solve a problem, you first need to know what it is you’re trying to solve. What’s the root of the problem?"
@@ F.3 Scheduler Pretraining
Specifically, we consider two settings. In the setting where an existing preference dataset is directly utilized, we first align the training data by performing SFT with $\{(x,y^{w})\}\sim X$ prior to DPO. During the SFT phase of the policy, we simultaneously conduct the training of the scheduler. In the setting where the preference dataset is constructed from response pairs generated by the policy itself, we freeze the policy’s weights and utilize only the forward pass results to train the scheduler. The algorithm for training the scheduler remains consistent with that described in Section 3. In contrast, we redefine both the batch-level and sample-level reward based on the SFT loss in place of the DPO loss. Specifically, we formally define the batch-level reward for round t1t-1:
@@
Among them, $y^{w}_{t,i,s}$ represents the $s$-th token of the $i$-th chosen response sequence at round $t$ .